The center of a circle is (1,2) and the point (3,13) is on the circle. What is the area of the circle?
Found 2 solutions by htmentor, Alan3354:
Answer by htmentor(1343)

The radius is the distance between the center (1,2) and the point (3,13)
The square of the radius, r^2 = (3-1)^2 + (13-2)^2 = 125
The area of the circle, A = pi*r^2 = pi*125 = 392.7

Answer by Alan3354(69443)

The center of a circle is (1,2) and the point (3,13) is on the circle. What is the area of the circle?
------------